# Introducing a Physics-informed Deep Learning Framework for Bridge Scour Prediction

### v2

@@ 3.2.3 Effect of Base DL Algorithm on SPINNs
NLinear 28,392 56,448
As shown in Table 7, there is a direct correlation between the model’s complexity and its computational cost. The CNN architecture requires approximately twice the computational effort of the LSTM due to its complex convolutional layers. In contrast, the NLinear model demands significantly fewer resources—only about 0.77% of the FLOPs required for LSTM operations. Regarding the SPINN framework, HEC18, TD, and GTD only introduce a few latent parameters to the base NN (See Table 4, thereby leading to a negligible extra computational cost. This enables SPINN variants to leverage the benefits of physics-informed learning without incurring a substantial computational burden.
Across the bridges analyzed in this research, the models founded on the NLinear consistently outperform the LSTM and CNN in terms of average MSE values. Considering the relatively small computational cost, NLinear works more efficiently compared to others. The potential reason behind this phenomenon has been unveiled by the authors of NLinear (Zeng et al., 2023). According to their experiments on various time series datasets, it has been found that linear models can be more effective in certain scenarios where there are clear trends and periodicity in the data. This is due to their simplicity and direct approach to handling time series data.
However, it is important to acknowledge that bridge-scouring events are a complex process. While NLinear models offer considerable advantages, a comprehensive modelling approach should not disregard the potential utility of other NN architectures. Different NN structures may provide unique insights or afford better performance under diverse conditions. When modelling a wider range of scenarios, multiple NN structures should be considered as baselines.
3.2.4 General DL Models for Bridge Clusters
In this study, we implemented Pure NN (All) and SPINN-GTD experimental groups to assess general DL models that can provide reasonable forecast accuracy across a cluster of bridges (such as our selected four case-study bridges in Alaska). General models can significantly reduce development and computational costs by eliminating the need for site-specific models (i.e., individual bridge models). As discussed in Section 3.2.1, the general data-driven models outperformed the site-specific data-driven models; however, as observed in Fig. 9, the general hybrid physics-data-driven models (SPINN-GTD) did not show superiority to site-specific physics-data-driven variants, i.e. SPINN-HEC18 (except for LSTM-539, LSTM-742, CNN-212) and SPINN-TD (except for CNN-212).
The challenge in generalization arises from the diverse riverbed, flow and pier geometry across different bridges. The GTD equation is designed to capture a broad range of scour behaviour within a single equation by combining multiple bridge-specific coefficients, i.e., $K_{1}$, $K_{2}$, $K_{3}$, and $a$ into one latent parameter $p_{3}$. When incorporating GTD as the physical loss, the equation could potentially limit the NN’s ability to adequately learn and adapt to the specific scouring patterns inherent to each bridge, leading to a degradation in prediction accuracy. Therefore, a more nuanced approach involving bridge-specific adaptations (such as coefficients used in HEC18 and TD), might be necessary to help improve NN learning.
@@ 3.3 Forecasting Trends and Capturing Scour Events
The primary objective of the scour forecasting model lies in its ability to predict major scour events in advance. Fig. 10 and Fig. 11 present the SPINN model prediction over the test dataset by the three SPINN models. Due to relatively stable bed elevation in the test sets of Bridge 212 and 527, our analysis here is focused on Bridges 539 and 742. The actual river bed elevation changes are depicted as a solid brown line, while the model predictions are shown as dashed lines. The predictions using pure NN models and the calibrated HEC18 empirical equation (CEE-SPINN-HEC18) derived from SPINN, as discussed in Section 3.4, are used as baselines for comparison. The magnified subplots in Fig. 10 and Fig. 11 highlight the analysis of major scour events on Bridges 539 and 742. The major scour event at bridges 539 and 742 occurs in July with a maximum depth of 1.5 and 1.4 meters, respectively. Table 8 compares the RMSE in maximum scour depth prediction for these two major events.
Table 8: Prediction error over the main scour events of Bridges 539 and 742. RMSEs ($m$) are computed between the starting and ending timesteps of the major souring events in their test sets. Bold values highlight the improved prediction of SPINN compared to Pure NN.
Model Training Datasets Method RMSE - 539 RMSE - 742
LSTM Site Specific Pure NN 0.506±0.057 0.289±0.011
LSTM Site Specific SPINN-HEC18 0.509±0.063 0.292±0.022
LSTM Site Specific SPINN-TD 0.514±0.025 0.264±0.012
LSTM General Pure NN 0.657±0.070 0.375±0.052
LSTM General SPINN-GTD 0.626±0.064 0.402±0.054
CNN Site Specific Pure NN 0.493±0.037 0.234±0.016
CNN Site Specific SPINN-HEC18 0.455±0.035 0.228±0.048
CNN Site Specific SPINN-TD 0.483±0.033 0.234±0.016
CNN General Pure NN 0.518±0.035 0.273±0.007
CNN General SPINN-GTD 0.525±0.026 0.259±0.015
NLinear Site Specific Pure NN 0.393±0.003 0.283±0.004
NLinear Site Specific SPINN-HEC18 0.396±0.003 0.284±0.007
NLinear Site Specific SPINN-TD 0.381±0.002 0.271±0.003
NLinear General Pure NN 0.488±0.004 0.243±0.002
NLinear General SPINN-GTD 0.490±0.003 0.243±0.001
Figure 10: Model predictions versus actual observations for Bridge 539 test set.
Figure 11: Model predictions versus actual observations for Bridge 742 test set.
The site-specific pure NN models show varying performance across the LSTM, CNN, and NLinear groups. The LSTM-based models tend to show more fluctuations in predicting scour episodes, while the CNN and NLinear groups generally align more closely with actual scour trends.
The performance of SPINN variants varies depending on the base model and the specific bridge. For Bridge 539, the SPINN-TD variant shows slightly less error in scour depth prediction for CNN and NLinear base models (2.03% and 3.05%, respectively). The SPINN-HEC18 variant also enhances the performance of the CNN model, reducing the RMSE by 7.71%. In the case of Bridge 742, SPINN-HEC18 and SPINN-TD groups also show similar improvement but are not consistent for all base models.
The general data-driven and SPINN models often show larger errors in capturing trends of scour and filling compared to site-specific models. Only for Bridge 742, the SPINN-GTD variant demonstrated improved performance compared to the site-specific pure NN models when paired with the CNN and NLinear base models.
@@
Comparing both data-driven and hybrid-physics-data-driven models with the calibrated HEC18 equation (CEE-SPINN-HEC18) shows the superiority of using dynamic DL models for scour prediction. This model struggles significantly in following the scouring and filling trends and significantly underestimating scour depth in most periods for both bridges 539 and 742 (see Figs. 10 and 11).
The performance of the models in capturing the major scour depth is generally consistent with their overall performance on the test dataset, as presented in Table LABEL:tab:table_overall_performance. Minor differences in the relative performance of the models between the main scouring event (Table 8) and the overall test set (Table LABEL:tab:table_overall_performance) can be attributed to the specific characteristics of the main scouring event and the model’s ability to capture those particular patterns.
Figure 12: The calibration of empirical equations with NLinear SPINN model, using the 539 dataset for site-specific models (GTD uses all datasets). a) The latent parameters of the HEC18 equation, b) The latent parameters of the TD equation, and c) The latent parameters of the GTD equation.
3.4 Calibrated Empirical Equations
@@
This section focuses towards the calibrated empirical equations derived from the SPINN framework. We will evaluate the performance of the three calibrated equations (see Table 8) and assess their potential to be implemented as independent scour models.
@@
Since the empirical equations are designed to predict maximum scour depth, their performance is only evaluated during scouring episodes. Bridge 539 paired with the NLinear group is chosen here as an example for detailed analysis, considering the deeper scouring depth and the superior performance of the base model.
@@
Fig. 12 displays the training progress plots of the latent parameters incorporated in the three empirical equations over 500 training epochs. The parameter $p_{1}$ in all equations acts as an overall adjustment factor, converging to distinct values for each equation, indicating its role in tailoring the output of each empirical equation. The $p_{2}$ in HEC18 and TD, representing the adjustment on the velocity, stabilizes at 1.0 after initial training epochs, suggesting its impact may be mitigated by adjustments in $p_{1}$. Similarly, $p_{3}$ in TD converges to 1.0, reflecting its minimal influence on the final results. $T_{L}$, the time lag of scour events, is calibrated to 1.95 and 2.35 hours in TD and GTD equations, respectively. This suggests the average time lag from start to stable scouring depth is around 2 hours for Bridge 539. Note that this length represents the average lag learned by the model, and it could be shorter or longer than particular cases. In the general model (GTD), the exponents $\alpha$ and $\beta$ display fluctuating convergence, indicating challenges in finding stable optimal values. The comprehensive summary of these calibrated parameters across different scenarios can be found in Table 9.
@@
212 HEC18 2.49 1.0 N/A N/A N/A N/A
212 TD 0.539 1.0 1.0 1.96 N/A N/A
527 HEC18 0.524 1.0 N/A N/A N/A N/A
527 TD 0.109 1.0 1.0 0.371 N/A N/A
539 HEC18 1.78 1.0 N/A N/A N/A N/A
539 TD 0.483 1.0 1.0 1.79 N/A N/A
742 HEC18 1.34 1.0 N/A N/A N/A N/A
742 TD 0.342 1.0 1.0 10.7 N/A N/A
All GTD 0.033 N/A N/A 2.35 0.987 0.697
Figure 13: a) Maximum scour prediction using the calibrated empirical models (CEE) for Bridge 539, b) magnified main scour episode.
Fig. 13 compares the calibrated empirical equations (CEE) derived from the SPINN model for predicting maximum scour depth on the Bridge 539 test dataset. The actual bed elevation changes are highlighted with a solid brown line, while the predictions from the CEEs are marked with dashed lines. The CEE-SPINN-TD and CEE-SPINN-GTD employ time-dependent equations, where an exponential function predicts the scour depth for each timestep in the scouring episode. The predicted depth is zero at the beginning of the episode ($t=0$) and reaches the maximum scour depth ($y_{s_{max}}$) at a time close to the calibrated time lag parameter ($T_{L}$). Given the multi-step nature of the CEE-SPINN-TD and CEE-SPINN-GTD, each timestep within the scouring episode is predicted multiple times based on its sequential positioning. The dashed lines represent the mean of these multiple predictions, while the shaded areas denote the 95% confidence intervals. On the other hand, the CEE-SPINN-HEC18 (New $E_{ref}$) uses the same dynamic reference level approach as the time-dependent equations to enable fair comparison. Since the CEE-SPINN-HEC18 (New $E_{ref}$) does not have multiple predictions for each timestep, it has a single dashed line without uncertainty shades.
Among the CEEs calibrated through SPINN, the HEC18 shows difficulty in following the trends of scouring events and overestimates the maximum scour depth by up to 3 meters. As expected, GTD’s performance is not as accurate as TD, which represents a site-specific version of the GTD model. During the main scouring event highlighted in Fig. 13b, the CEE-SPINN-HEC18, CEE-SPINN-TD, and CEE-SPINN-GTD yield RMSEs ($m$) of 2.608, 0.562, and 1.096, respectively. While not as accurate as the SPINN models, these results demonstrate that the SPINN-calibrated time-dependent empirical equation can provide reasonable estimates of maximum scour depth with the advantage of much greater simplicity compared to the neural networks.
@@ 4   Conclusion
This paper introduced scour physics-informed neural network algorithms, SPINNs, exploring hybrid physics-data-driven models for bridge scour prediction based on deep learning. Central to SPINN’s methodology is the integration of physics-based empirical equations as supplementary loss components into deep learning neural networks. We focused on three deep learning architectures, LSTM, CNN, and NLinear, as base models for scour prediction. Both CNN and LSTM models have already proved competitive in real-time scour forecasts using historical monitoring data in our previous studies (Yousefpour and Correa, 2023; Hashem and Yousefpour, 2024). NLinear model was introduced for the first time for scour forecasting, demonstrating the highest accuracy and much lower computational cost compared to CNN and LSTM. Although SPINNs performance showed variability with the base DL model, the physics-based loss function/empirical equation, and the bridge-specific data, overall they showed superior performance in most forecasting scenarios/case studies compared to pure data-driven NN models. For some scenarios, such as the LSTM group in Bridge 539, the SPINN could help the base model reduce up to 50% forecasting error.
We explored the idea of generalised models for bridge clusters by aggregating bridge datasets in Alaska to train the DL models and use them for each specific bridge. The pure NN models mostly benefited from this approach. In particular, bridges with limited datasets achieved much more accuracy in predictions with the general models compared to site-specific ones. However, the general hybrid physics-data-driven models (SPINN-GTD) did not show superiority to site-specific physics-data-driven variants. We found that bridge-specific variants of SPINN, incorporating HEC18 and time-dependent (TD) empirical equations yield more accurate predictions compared to general time-dependent (GTD). The GTD empirical equation needs to adapt to a wide range of scour behaviour for a cluster of bridges during training, i.e., finding an optimal solution for latent parameters that are, in fact, bridge-specific. This can inadvertently limit the NN’s ability to adequately learn and adapt to the specific scouring patterns in each particular bridge.
This study revealed that the sister calibrated empirical equations derived from SPINN, particularly the introduced time-dependent HEC18, can provide reasonable estimates of the maximum scour depth with potentially much higher accuracy than the traditional methods, such as HEC-18. As opposed to traditional HEC-18, which is often calibrated using discrete scour measurements, the empirical equations within SPINN are calibrated using historical time series scour monitoring data throughout the neural network training process. The generalised time-dependant calibrated equations can be particularly helpful in scenarios where site-specific bridge, flow and geology data are lacking, such as new bridge projects in a region.
Although the error reduction margins in scour predictions varied across different baseline models and bridges, SPINN framework proved to be an instrumental approach to improve scour prediction both in real-time forecasting and maximum scour depth estimation scenarios. Comparing the HEC18 model with SPINNs and pure DL models shows a significant improvement, which can benefit the current state of practice in bridge scour design and management.
@@ 6   Acknowledgments
212 NLinear Pure NN N/A 212 0.010±0.001
212 NLinear Pure NN (General) N/A All 0.013±0.000
212 NLinear SPINN HEC18 212 0.010±0.000
212 NLinear SPINN TD 212 0.010±0.000
212 NLinear SPINN GTD All 0.013±0.000
@@
527 NLinear Pure NN N/A 527 0.019±0.000
527 NLinear Pure NN (General) N/A All 0.049±0.001
527 NLinear SPINN HEC18 527 0.019±0.000
527 NLinear SPINN TD 527 0.023±0.000
527 NLinear SPINN GTD All 0.049±0.001
@@
539 NLinear Pure NN N/A 539 0.105±0.000
539 NLinear Pure NN (General) N/A All 0.168±0.002
539 NLinear SPINN HEC18 539 0.105±0.002
539 NLinear SPINN TD 539 0.114±0.001
539 NLinear SPINN GTD All 0.169±0.001
@@
742 NLinear Pure NN N/A 742 0.199±0.001
742 NLinear Pure NN (General) N/A All 0.310±0.004
742 NLinear SPINN HEC18 742 0.200±0.002
742 NLinear SPINN TD 742 0.229±0.001
742 NLinear SPINN GTD All 0.310±0.002

### v3

@@ 3.6 Evaluation of Base DL Algorithms for SPINNs
As shown in Table 8, there is a direct correlation between the model’s complexity and its computational cost. The CNN architecture requires approximately twice the computational effort of the LSTM due to its complex convolutional layers. Regarding the SPINN framework, HEC18, TD, and GTD only introduce a few latent parameters to the base NN (See Table 3), thereby leading to a negligible extra computational cost. This enables SPINN variants to leverage the benefits of physics-inspired learning without incurring a substantial computational burden.
LSTM shows to achieve significant performance improvement in terms of accuracy (error) and robustness (variance) from a physics-inspired architecture. Given the variability of scour data in various bridges, it is important to investigate competing NN base models to achieve the most optimal performance.
Figure 9: Model predictions versus actual observations for Bridge 539 test set (May-September 2021). Top graph displays the flow changes in this period. (a-c) show LSTM-based predictions, (d-f) display CNN-based predictions, and each row represents SPINN-HEC18, SPINN-TD, and SPINN-GTD, respectively. Additional test periods are presented in the Appendix.
Figure 10: Model predictions versus actual observations for Bridge 742 test set (July-September 2017). Top graph displays the flow changes in this period. (a-c) show LSTM-based predictions, (d-f) display CNN-based predictions, and each row represents SPINN-HEC18, SPINN-TD, and SPINN-GTD, respectively.
@@ 3.7 Forecasting Trends and Capturing Scour Events
The primary objective of the scour forecasting model lies in its ability to predict major scour and filling events in advance. Fig. 9 and Fig. 10 present the SPINN model predictions over the test dataset for the three SPINN models. Due to relatively stable bed elevation in the test sets of Bridge 212 and 527, our analysis here is focused on Bridges 539 and 742. The actual river bed elevation changes are depicted as a solid brown line, while the model predictions are shown as dashed lines. The predictions using pure NN models and the calibrated HEC18 empirical equation (CEE-SPINN-HEC18) derived from SPINN (see Section 3.8) are used as baselines for comparison (see also Table 6).
@@
Comparing both data-driven and hybrid-physics-data-driven models with the calibrated HEC18 equation (CEE-SPINN-HEC18) shows the superiority of the DL models for scour prediction compared with an empirical model. The calibrated HEC-18 model struggles to follow the scouring and filling trends and significantly underestimates the scour depth in most periods for bridges 539 and 742 (see Figs. 9 and 10).
3.8 The SPINN-Calibrated Empirical Equations
@@
In this section, we shift the focus towards the calibrated empirical equations derived from the proposed SPINN framework. We will evaluate the performance of the three calibrated equations (see Table 3) and assess their potential to be implemented as independent scour models.
@@
Since the empirical equations are designed to predict maximum scour depth, their performance is only evaluated during scouring episodes. Bridge 539 paired with the LSTM group is chosen here as an example for detailed analysis, considering the deeper scouring depth and the superior performance of the base model.
Figure 11: The calibration of empirical equations with LSTM SPINN model, using the 539 dataset for site-specific models (GTD uses all datasets). a) The latent parameters of the HEC18 equation, b) The latent parameters of the TD equation, and c) The latent parameters of the GTD equation.
@@
Fig. 11 displays the training progress plots of the latent parameters incorporated in the three empirical equations over 500 training epochs. The parameter $p_{1}$ in all equations acts as an overall adjustment factor, converging to distinct values for each equation, indicating its role in tailoring the output of each empirical equation. The $p_{2}$ in HEC18 and TD, representing the adjustment on the velocity, stabilizes at 1.0 after initial training epochs, suggesting its impact may be mitigated by adjustments in $p_{1}$. Similarly, $p_{3}$ in TD converges to 1.0, reflecting its minimal influence on the final results. $T_{L}$ is calibrated to various values for site-specific models and to 2.3 hrs for the general model. The summary of calibrated parameters across different scenarios can be found in Table 9.
@@
212 HEC18 2.480 1.000 N/A N/A N/A N/A
212 TD 0.534 1.000 1.000 1.970 N/A N/A
527 HEC18 0.371 1.000 N/A N/A N/A N/A
527 TD 0.098 1.000 1.000 0.356 N/A N/A
539 HEC18 1.790 1.000 N/A N/A N/A N/A
539 TD 0.514 1.000 1.000 1.750 N/A N/A
742 HEC18 1.350 1.000 N/A N/A N/A N/A
742 TD 0.329 1.000 1.000 9.470 N/A N/A
All GTD 0.076 N/A N/A 2.290 0.903 0.592
Figure 12: Maximum scour prediction using the calibrated empirical models (CEE) for Bridge 539, a) the overview of bed elevation changes; (b-e) prediction on each scour episode using three SPINN-calibrated equations.
Fig. 12 compares the calibrated empirical equations (CEE) derived from the SPINN model for predicting maximum scour depth on the Bridge 539 test dataset. The actual bed elevation changes are highlighted with a solid brown line, while the predictions from the CEEs are marked with dashed lines. The CEE-SPINN-TD and CEE-SPINN-GTD employ time-dependent equations, where an exponential function predicts the scour depth for each timestep in the scouring episode. The predicted depth is zero at the beginning of the episode ($t=0$) and reaches the maximum scour depth ($y_{s_{max}}$) at a time close to the calibrated time lag parameter ($T_{L}$). On the other hand, the CEE-SPINN-HEC18 uses the same dynamic reference level approach as the time-dependent equations to estimate the maximum scour depth of each episode (horizontal dashed line).
Among the CEEs calibrated through SPINN, the HEC18 overestimates the maximum scour depth by up to 3 meters. Across these scouring episodes in Fig. 12, the CEE-SPINN-HEC18, CEE-SPINN-TD, and CEE-SPINN-GTD yield RMSEs ($m$) of 2.539, 0.443, and 0.34 respectively. Although not as accurate as SPINN models, the proposed SPINN-calibrated time-dependent empirical equations show great potential to replace existing empirical models for maximum scour depth estimation.
@@ 4   Conclusion
This paper introduced scour physics-inspired neural network algorithms, SPINNs, exploring hybrid physics-data-driven models for bridge scour prediction based on deep learning. Central to SPINN’s methodology is the integration of physics-based empirical equations as supplementary loss components into deep neural networks. Although SPINNs performance showed variability with the choice of base DL model, the empirical equation and the case study bridge, they showed superior performance in the majority of scenarios compared to pure data-driven (NN) models. The bridge-specific variants of SPINN with the proposed time-dependent (TD) empirical equation showed the most competitive performance among other SPINN variants. For some scenarios, such as the LSTM group for Bridge 539, the SPINN could help the base NN model reduce forecasting error by up to 50%.
We explored the idea of transferable models for bridge clusters by aggregating bridge datasets across a cluster to train general DL models that can provide reasonable scour prediction for any bridge within that cluster. The general pure data-driven (NN) models showed superiority over site-specific models, however, the general SPINN with the proposed empirical model did not provide as accurate predictions as site-specific SPINNs.
This study revealed that the sister calibrated empirical equations derived from SPINN framework, particularly the introduced time-dependent equation, can provide a viable alternative to traditional scour models, such as FHWA HEC-18. As opposed to traditional HEC-18, which is often calibrated using discrete scour lab measurements, the empirical equations within SPINN are calibrated using historical time series of scour monitoring data throughout the neural network training process. The proposed general time-dependent calibrated equation can be particularly helpful in scenarios where site-specific flow and riverbed conditions are not readily available (e.g., in new bridge projects).
Comparing SPINNs and the SPINN-calibrated empirical equations with traditional empirical models presents a significant improvement, which can benefit the current state of practice in bridge scour design and management. This paper can pave the path for future studies on physics-inspired machine learning for scour prediction.
@@ 6   Acknowledgments
Figure A4: Model predictions versus actual observations for Bridge 539 test set (July-October 2019). Top graph displays the flow changes in this period. (a-c) show LSTM-based predictions, (d-f) display CNN-based predictions, and each row represents SPINN-HEC18, SPINN-TD, and SPINN-GTD, respectively.
Figure A5: Model predictions versus actual observations for Bridge 539 test set (July-September 2020). Top graph displays the flow changes in this period. (a-c) show LSTM-based predictions, (d-f) display CNN-based predictions, and each row represents SPINN-HEC18, SPINN-TD, and SPINN-GTD, respectively.